OpenStudy (anonymous):

Assuming x > 0, what is the simplified fraction form for 15 x times the square root of 3, all divided by the square root of 75 x squared.

OpenStudy (zehanz):

I think you have this: \[\frac{ 15x \sqrt3 }{ \sqrt{75x^2} }\]I would start with simplifying the denominator: square roots can be easily calculated as long as you have squares ;) So try to get as much squares as you can:\[\sqrt{75x^2}=\sqrt{25x^2 \cdot 3}=\sqrt{25x^2} \cdot \sqrt{3}=...\]Can you do this?

OpenStudy (anonymous):

so then I would have \[5x \times \sqrt{3}\]?

OpenStudy (zehanz):

Yes, so now it is:\[\frac{ 15x \sqrt{3} }{ 5x \sqrt{3} }=\]Well, after simplifying this fraction, there won't be much left...